Abu ran for 1/3 of an hour and then walked for half an hour at the rate of 3 km/h. In all, he had covered 4 km. How fast did he run?
Found 2 solutions by mananth, Maynard:
Answer by mananth(16949)

let Abu's running rate be x km/h
he ran for 1/3 hour
so he ran a distance = 1/3 * x= x/3 km
..
he walked for 1/2 hour speed = 3 km/h
d= 1/2 * 3 = 3/2 km
..
walking distance + running distance = 4 km
3/2 +x/3 =4
-3/2 -3/2
x/3 = 4-3/2
x/3 = (8-3)/2
x/3 = 5/2
x = 3*5 /2
x= 15/2
..
running speed = 7.5 km/h
...
m.ananth@hotmail.ca

Answer by Maynard(1)